b1az3:

Find the function h(x) = f(x) + g(x) if f(x) = 4x + 2 and g(x) = x - 4. h( x) = 4 x + 1 -2 h( x) = 4 x + x + 6 h( x) = 4 x - 4 h( x) = 4 x + x - 2

b1az3:

h( x) = 4^x + x - 2

AZ:

just add h(x) = f(x) + g(x) f(x) = 4x + 2 g(x) = x - 4

AZ:

yup, that's correct!

AZ:

oh there's an exponent

AZ:

well, it's still correct